OpenStudy (grant330sims):

simplify. 8x-56/x^2-49 times x^2+11x+28/x-6

OpenStudy (anonymous):

very simple...

OpenStudy (anonymous):

8x-56/x^2-49 =8(x-7) / (x-7) (x+7) =8/ (x+7)

OpenStudy (anonymous):

x^2+11x+28/x-6 =(x+7)(x+4) / (x-6)

OpenStudy (anonymous):

understand?

OpenStudy (grant330sims):

so the final answer is 8(x+4)/x-6?